3. You're trying to determine whether to expand your business by building a new manufacturing plant. The plant has an installation cost of $12 million, which will be depreciated straight-line to zero over its 4-year life. The plant has projected net income of $1,095,000, $902,000, $1,412,000, and $1,724,000 over these 4 years. What is the average accounting return?
Answer #1

Average net income=Total net income/Total time period

=(1,095,000+902,000+1,412,000+1,724,000)/4=$1283250

Average investment=(12,000,000/2)=$6,000,000

AAR=Average net income/Average investment

=1283250/6,000,000

=21.3875%